### v1

@@ Abstract
In this paper, based on Birkhoff orthogonality, we introduce two geometric constants $A_{t}^{B}(X)$ and $D_{t}^{B}(X)$ in Banach spaces, which generalize the skew geometric constants related to Birkhoff orthogonality. We systematically investigate the basic properties of the two constants, including their upper and lower bounds, and establish the equivalent characterizations for Banach spaces being uniformly non-square. Additionally, we explore the relationship between $D_{t}^{B}(X)$ and the modulus of convexity $\delta_{X}(\varepsilon).$ Finally, we explore several applications of the two newly proposed geometric constants.
@@ 1 Introduction
Ji D. and Wu S. introduced the constants $D(X)$ and $D^{\prime}(X)$ in [28] to measure the distinction between Birkhoff orthogonality and isosceles orthogonality.
@@
Suppose $X$ is a Banach space, for every $\varepsilon\in[0,2]$, we define the modulus of convexity of the norm $\|\cdot\|$ to be ([23])
@@
In Section 3, we consider the $D_{t}^{B}(X)$, a new geometric constant. Then we present several basic properties of this constant, including its upper and lower bounds as well as its connection to the uniform non-squareness of Banach spaces. We also establish the relation between $D_{t}^{B}(X)$ and the modulus of convexity $\delta_{X}(\varepsilon)$.
@@ 2 The constant $A^{B}_{t}(X)$
We define the constant $A^{B}_{t}(X)$ by extending and generalizing the form of the existing Birkhoff orthogonality-related geometric constant $A_{2}(X,B)$ by introducing a parameter $t>0$. We suppose that $X$ is a Banach space with dimension at least 2. Next we start by outlining the key definitions: for $t>0$,
@@
$A^{B}_{1}(X)$ $=A^{B}_{t}(X)$
$=\sup\left\{\frac{\|x+y\|+\|x-y\|}{2}:x,y\in S_{X},x\perp_{B}y\right\}=A_{2}(X,B).$
@@
$\frac{t\left(\left\|\frac{1}{t}x+y\right\|+\left\|x-\frac{1}{t}y\right\|\right)}{2t}\geqslant\frac{\|x+y\|+\|x-y\|}{2}.$
$\frac{A^{B}_{t}(X)}{t}\geqslant A_{2}(X,B).$
In short, we have
$A^{B}_{t}(X)\geqslant\min\{1,t\}A_{2}(X,B),$
On the other hand, for the above of $x,y$, by the Lemma 2, we have
@@
$\|tx-y\|\leq\|tx\|+\|y\|-\left(2-\left\|\frac{tx}{\|tx\|}-\frac{y}{\|y\|}\right\|\right)\min(\|tx\|,\|y\|)$
@@
By observate Proposition 1 and Proposition 3, when $A^{B}_{t}(X)=1+t$ we deduce that
@@
This shows that $A_{2}(X,B)\geq 2$. It should be noted that the upper bound of the constantis 2, and combined with the conclusion that $A_{2}(X,B)=2$ if and only if the space $X$ is not uniformly nonsquare([5]), as desired. ∎
@@ 3 The constant $D_{t}^{B}(X)$
In this part, we first introduce an orthogonality concept that is an extension of isosceles orthogonality.
@@
txytx-$y$-ytxtxxxty+xty+xyy
Figure 1: An example of skew isosceles orthogonality.
@@
Based on the constants studied previously, we introduce a new constant $D_{t}^{B}(X)$ and investigate some of its properties. The constant incorporates the parameter $t>0$, which is used as the denominator for normalization, ensuring that the value of the constant is independent of the magnitude of $t$. The new constant characterizes the discrepancy between Birkhoff orthogonality and skew isosceles orthogonality.
Suppose that $X$ is a real Banach space with dimension at least 2. We first present the key definitions as follows: for $t>0$,
@@
Let $X$ be a Banach space with $\operatorname{dim}X\geq 3$. Then
@@
According to the results on Birkhoff orthogonality in [35], there exist unit vectors $x,y\in S_{X}$ such that erxcit $x\perp_{B}y$ and $y\perp_{B}x$. Consider the two pairs $(x,y)$ and (y,x)(-y,x), we can obtain
@@
On the other hand, for any $x,y\in S_{X}$, $t>0$, and $x\perp_{B}y$, we can obtain that
@@
we can proof that $D_{t}^{B}(X)\leq\frac{1}{t}.$
When the dimension of the space is sufficiently large, the lower bound can be explained in an alternative way, thanks to the celebrated Dvoretzky’s theorem.
@@
Suppose that $X$ is an inner product space, by the Lemma 5, we have
$\|x+ty\|-\|tx-y\|=\|x+ty\|-\|tx+y\|=0,$
for all $t>0$ and $x,y\in S_{X}$, hence we have $D_{t}^{B}(X)=0$. ∎
@@
Let $X=\mathbb{R}^{2},0<t\leq 1$ and assign the following $\ell_{\infty}-\ell_{1}$ norm
@@
Then $D_{t}^{B}(X)=1$.
Suppose that x=(0,1)x=(0,-1), $y=(1,0)$, it is easy to see that $x,y\in S_{X}$ and $x\perp_{B}y$. We can direct calculation gives $\|x+ty\|=1+t$ and $\|tx-y\|=1$. So we can obtain that
$\frac{\|x+ty\|-\|tx-y\|}{t}=\frac{1+t-1}{t}=1.$
@@
The upper bound $\frac{1}{t}$ of $D_{t}^{B}(X)$ is attained by a pair of points of $S_{X}$ if and only if there exist two points $x,y\in S_{X}$ such that $[x,y]$ and $\left[x,x-\frac{2}{t}y\right]$ are both contained in $S_{X}$. Correspondingly, the length of the segment $[x,y]$ is at least $\frac{1}{t}$.
@@
This means that the midpoint of $x$ and $\left[x,x-\dfrac{2}{t}y\right]$ has norm 1, then we have $\left[x,x-\dfrac{2}{t}y\right]\subset S_{X}$. In addition, since $x\perp_{B}y$, we have $\|x-y\|\geq 1$.
@@
Proposition 10.
Let $X$ be Banach space, if $D_{t}^{B}(X)=1$ then $\delta_{X}(1)=0.$
@@
If $D_{t}^{B}(X)=1$, we assume that $p\in\mathbb{N}$, then we have two points $x_{p},y_{p}\in S_{X}$ satisfying $x_{p}\perp_{B}y_{p}$ and $t_{p}>0$ such that
$\frac{\left\|x_{p}+t_{p}y_{p}\right\|-\left\|t_{p}x_{p}-y_{p}\right\|}{t_{p}}>1-\frac{1}{p}.$
@@
$t_{p}\left\|x_{p}+y_{p}\right\|$ $\geq\left\|x_{p}+t_{p}y_{p}\right\|-\left\|\left(1-t_{p}\right)x_{p}\right\|$
$>\left(1-\frac{1}{p}\right)t_{p}+\left\|t_{p}x_{p}-y_{p}\right\|-1+t_{p}$
$\geq\left(1-\frac{1}{p}\right)t_{p}+1-1+t_{p}$
$=\left(2-\frac{1}{p}\right)t_{p}.$
Hence $\left\|x_{p}+y_{p}\right\|>2-\frac{1}{p}$ . Therefore we have
@@
$\left\|x_{p}+y_{p}\right\|$ $\geq\left\|x_{p}+t_{p}y_{p}\right\|-\left\|\left(t_{p}-1\right)y_{p}\right\|$
$>\left(1-\frac{1}{p}\right)t_{p}+\left\|t_{p}x_{p}-y_{p}\right\|-1+t_{p}$
$\geq\left(1-\frac{1}{p}\right)t_{p}+1-1+t_{p}$
$=\left(2-\frac{1}{p}\right)t_{p}.$
@@
Suppose that $x,y\in S_{X}$ and $x\perp_{B}y$, then we have $\|tx-y\|\geq t$, By the definition of $\delta_{x}(1)$, we can obtain that $\|x+y\|\leq 2\left(1-\delta_{x}(1)\right)$.
@@
We will establish the respective connections among $J^{B}(X)$, $D^{\prime}(X)$ and $D_{t}^{B}(X)$ in the subsequent sections.
@@
Note that a Banach space $X$ is uniformly non-square if and only if $J^{B}(X)<2$ [12], by the above Proposition 12, we deduce the following Proposition.
Proposition 13.
@@
This shows that $J^{B}(X)\geq 2$. It should be noted that the upper bound is 2, and combined with the conclusion that $J^{B}(X)=2$ if and only if the space $X$ is not uniformly nonsquare, as desired. ∎
Theorem 14.
Let $X$ be Banach space, then
$\frac{(1+t)D^{\prime}(X)-2|t-1|}{2t}\leq D_{t}^{B}(X)\leq\frac{tD^{\prime}(X)+2|1-t|}{t}.$
Proof.
On the one hand,
$\frac{\|x+ty\|-\|tx-y\|}{t}$ $\leq\frac{(t\|x+y\|+|1-t|)-(t\|x-y\|-|1-t|)}{t}$
$=\frac{t(\|x+y\|-\|x-y\|)+2|1-t|}{t}.$
So we can obtain,
$\frac{\|x+ty\|-\|tx-y\|}{t}\leq\frac{tD^{\prime}(X)+2|1-t|}{t}.$
On the other hand, we have
$\frac{\|x+ty\|-\|tx-y\|}{t}$
$\geq\frac{\left(\left\|\frac{1+t}{2}(x+y)\right\|-\left\|\frac{1-t}{2}(x-y)\right\|\right)-(\|x-y\|+|t-1|\|x\|)}{t}$
$=\frac{\frac{1+t}{2}\|x+y\|-\frac{1+t}{2}\|x-y\|-|t-1|}{t}.$
Hence,
$D_{t}^{B}(X)\geq\frac{1+t}{2t}D^{\prime}(X)-\frac{|t-1|}{t}.$
So we have
$\frac{(1+t)D^{\prime}(X)-2|t-1|}{2t}\leq D_{t}^{B}(X)\leq\frac{tD^{\prime}(X)+2|1-t|}{t},$
as desired. ∎
@@ 4 Some applications of $A^{B}_{t}(X)$ and $D_{t}^{B}(X)$
Recall that an orthogonality relation $\perp$ is calls symmetric if the condition $x\perp y$ necessarily entails $\perp x$ . It is straightforward to verify that the canonical orthogonality defined on Hilbert spaces is symmetric. In contrast, it is a well-documented fact that Birkhoff orthogonality lacks this symmetry property. Nevertheless, James established the following conclusion in the work ([24]).
@@
([24]) A Banach space $X$ with dimension $\operatorname{dim}X\geq 3$ is a Hilbert space if and only if Birkhoff orthogonality is symmetric on $X$.
@@
([25]) A two-dimensional Banach space in the Birkhoff orthogonality is symmetric, defined as a Radon plane.
@@
From Proposition 3 and the above statement (1), the following conclusion holds.
Corollary 1.
p+qp+qppq-qpq-p-qp-pqq
Figure 3: The unit sphere of the Affine regular hexagon
By an affine-regular hexagon, any non-degenerate affine image obtained from a regular hexagon via affine transformation([26]). We observe that the space $\mathbb{R}^{2}$ with the norm $\ell_{\infty}-\ell_{1}$ is a Radon plane ([24]) (see Figure 2) and its unit sphere is an affine-regular hexagon. we have below example to achieve the upper bound of the corollary.
Example 3.
Then $A^{B}_{t}(X)=1+t-\frac{1}{2}\min\{1,t\}.$
Proof.
Suppose that $x=(1,0)$ and $y=(0,1)$. We can infer that $x,y\in S_{X}$ such that $x\perp_{B}y$ from Figure 1. In addition, when $t\geq 1$, we have $\|x+ty\|=1$, $\|tx-y\|=t+1.$ Thus, it follows from Corollary 1 that we derive
$1+\frac{t}{2}=\frac{\|x+ty\|+\|tx-y\|}{2}\leq A^{B}_{t}(X)=1+\frac{t}{2},$
which implies that $A^{B}_{t}(X)=1+\frac{t}{2}.$
Similary, when $0\leq t\leq 1$, we have $\|x+ty\|=t$, $\|tx-y\|=t+1,$ so we can obtain that $A^{B}_{t}(X)=t+\frac{1}{2}.$
Thus we can complete the example.
Theorem 15.
Proof.
Suppose that an affine-regular hexagon is $S_{X}$ , we have distinct points $p,q\in S_{X}$ such that $\pm p,\pm(p+q),\pm q$ are the vertices of $S_{X}$ (see Figure 3). Take $x=p+q$ and $y=p.$ Then it is evident that $x,y\in S_{X}$ with $x\perp_{B}y$.
In addition, we are also able to obtain
$\|tx-y\|=\|tq-(1-t)p\|=\|tq+(1-t)(-p)\|=1.$
Hence, by the definition of $A^{B}_{t}(X)$, we have
$1+\frac{t}{2}=\frac{\|x+ty\|+\|tx-y\|}{2}\leq A^{B}_{t}(X)\leq 1+\frac{t}{2},$
which means that $A^{B}_{t}(X)=1+\frac{t}{2}$.
Similary, $A^{B}_{t}(X)=t+\frac{1}{2}$, this completes the proof. ∎

### v2

@@ Abstract
In this paper, based on Birkhoff orthogonality, we introduce two geometric constants $A_{t}^{B}(X)$ and $D_{t}^{B}(X)$ in Banach spaces, which generalize the existing geometric constants. Moreover, one of these constants is investigated in connection with the newly introduced orthogonality relation. We systematically study the basic properties of the two constants, including their upper and lower bounds, and establish equivalent characterizations for a Banach space to be uniformly non-square. Additionally, we explore the relationship between $D_{t}^{B}(X)$ and the modulus of convexity. Finally, we present several applications of the two newly proposed geometric constants.
@@ 1 Introduction
Ji and Wu introduced the constants $D(X)$ and $D^{\prime}(X)$ in [28] to measure the distinction between Birkhoff orthogonality and isosceles orthogonality.
@@
Suppose $X$ is a Banach space, for every $\varepsilon\in[0,2]$, we define the modulus of convexity of the norm $\|\cdot\|$ to be [37]
@@
In Section 3, we consider the $D_{t}^{B}(X)$, a new geometric constant. Then we present several basic properties of this constant, including its upper and lower bounds as well as its connection to the uniform non-squareness of Banach spaces. We also establish the relation between $D_{t}^{B}(X)$ and the modulus of convexity.
@@ 2 The constant $A^{B}_{t}(X)$
We define the constant $A^{B}_{t}(X)$ by extending and generalizing the form of the existing Birkhoff orthogonality-related geometric constant $A_{2}(X,B)$ by introducing a parameter $t>0$. We suppose that $X$ is a Banach space with dimension at least 2. Next we start by outlining the key definitions: for fixed $t>0$,
@@
$A^{B}_{1}(X)=\sup\left\{\frac{\|x+y\|+\|x-y\|}{2}:x,y\in S_{X},x\perp_{B}y\right\}=A_{2}(X,B).$
@@
$\frac{t\left(\left\|\frac{1}{t}x+y\right\|+\left\|x-\frac{1}{t}y\right\|\right)}{2t}\geqslant\frac{\|x+y\|+\|x-y\|}{2},$
which implies that $\frac{A^{B}_{t}(X)}{t}\geqslant A_{2}(X,B)$.
On the other hand, by the Lemma 2, we have
@@
$\|tx-y\|\leq\|tx\|+\|y\|-\left(2-\left\|\frac{tx}{\|tx\|}-\frac{y}{\|y\|}\right\|\right)\min(\|tx\|,\|y\|).$
@@
By Proposition 3, when $A^{B}_{t}(X)=1+t$ we deduce that
@@
This shows that $A_{2}(X,B)\geq 2$. It should be noted that the upper bound of the constant is 2, and combined with the conclusion that $A_{2}(X,B)=2$ if and only if the space $X$ is not uniformly [5], as desired. ∎
@@ 3 The constant $D_{t}^{B}(X)$
Lemma 6, which will be used in this chapter, establishes a result concerning skew equality that is both meaningful and interesting. In particular, it can be used to characterize inner product spaces. Next, we introduce an orthogonality notion extending isosceles orthogonality, which is also skew.
@@
txytx-$y$-ytxtxxxx+tyx+tyyy
Figure 1: Illustration of skew isosceles orthogonality
@@
Based on the constants studied previously and the definition of skew isosceles orthogonality, we introduce a new constant $D_{t}^{B}(X)$ and investigate some of its properties. This new constant characterizes the discrepancy between Birkhoff orthogonality and skew isosceles orthogonality.
Suppose that $X$ is a Banach space with dimension at least 2. We first present the key definitions as follows: for fixed $t>0$,
@@
Remark 1.
The new constant $D_{t}^{B}(X)$ is also a generalization of $D^{\prime}(X)$ .
Lemma 4.
Let $X$ be a Banach space. Then there exist unit vectors $x,y\in S_{X}$ such that $x\perp_{B}y$ and $y\perp_{B}x$.
Proof.
According to the results on Birkhoff orthogonality in [35], there exist unit vectors $x_{0},y_{0}\in X$ such that $x_{0}\perp_{B}y_{0}$ and $y_{0}\perp_{B}x_{0}$. Combined with the homogeneity of Birkhoff orthogonality, it follows that there exist $x,y\in S_{X}$ such that $x\perp_{B}y$ and $y\perp_{B}x$.
@@
Let $X$ be a Banach space. Then
@@
From Lemma 4, there exist unit vectors $x,y\in S_{X}$ such that $x\perp_{B}y$ and $y\perp_{B}x$. Consider the two pairs $(x,y)$ and (y,x)(-y,x), we can obtain
@@
On the other hand, for any $x,y\in S_{X}$ and $x\perp_{B}y$, we can obtain that
@@
as desired.
When the dimension of the space is sufficiently large, the lower bound admits an alternative interpretation by virtue of the celebrated Dvoretzky’s theorem, a profound result in functional analysis. The values of geometric constants are determined by the structure of the space, and this theorem characterizes the subspace structure of high-dimensional Banach spaces. For studies on geometric constants via Dvoretzky’s theorem, we refer the reader to [29], [38] .
@@
Suppose that $X$ is an inner product space. By Lemma 6, combined with the fact that Roberts orthogonality and Birkhoff orthogonality are equivalent in inner product spaces, we have
$\|x+ty\|-\|tx-y\|=\|x+ty\|-\|x-ty\|=0$
for $x,y\in S_{X}$, and hence $D_{t}^{B}(X)=0$.
@@
Let $X=\mathbb{R}^{2},t\geq 1$ and assign the following $\ell_{\infty}-\ell_{1}$ norm
@@
Then $D_{t}^{B}(X)=\frac{1}{t}$.
Suppose that $x=(1,0)$, $y=(1,1)$, it is easy to see that $x,y\in S_{X}$ and $x\perp_{B}y$. We can direct calculation gives $\|x+ty\|=1+t$ and $\|tx-y\|=t$. So we can obtain that
$\frac{\|x+ty\|-\|tx-y\|}{t}=\frac{1+t-t}{t}=\frac{1}{t}.$
@@
The upper bound $\frac{1}{t}$ of $D_{t}^{B}(X)$ is attained by a pair of points of $S_{X}$ if and only if there exist two points $x,y\in S_{X}$ such that $[x,y]$ and $\left[x,x-\frac{2}{t}y\right]$ are both contained in $S_{X}$. Correspondingly, the length of the segment $[x,y]$ is at least $1$.
@@
This means that the midpoint of $x$ and $x-\dfrac{2}{t}y$ has norm 1, then we have $\left[x,x-\dfrac{2}{t}y\right]\subset S_{X}$. In addition, since $x\perp_{B}y$, we have $\|x-y\|\geq 1$.
@@
Theorem 10.
Let $X$ be Banach space, if $D_{t}^{B}(X)=\frac{1}{t}$ then $\delta_{X}(1)=0.$
@@
If $D_{t}^{B}(X)=\frac{1}{t}$, we assume that $p\in\mathbb{N}$, then we have two points $x_{p},y_{p}\in S_{X}$ satisfying $x_{p}\perp_{B}y_{p}$ and $t_{p}>0$ such that
$\frac{\left\|x_{p}+t_{p}y_{p}\right\|-\left\|t_{p}x_{p}-y_{p}\right\|}{t_{p}}>\frac{1}{t}-\frac{1}{p}.$
@@
$t_{p}\left\|x_{p}+y_{p}\right\|$ $\geq\left\|x_{p}+t_{p}y_{p}\right\|-\left(1-t_{p}\right)\left\|x_{p}\right\|$
$>\left(\frac{1}{t}-\frac{1}{p}\right)t_{p}+\left\|t_{p}x_{p}-y_{p}\right\|-1+t_{p}$
$\geq\left(\frac{1}{t}-\frac{1}{p}\right)t_{p}+t_{p}-1+t_{p}$
$=\left(\frac{1}{t}-\frac{1}{p}+2\right)t_{p}-1.$
Hence $\left\|x_{p}+y_{p}\right\|>\frac{1}{t}-\frac{1}{p}+2-\frac{1}{t_{p}}$ . Therefore we have
@@
$\left\|x_{p}+y_{p}\right\|$ $\geq\left\|x_{p}+t_{p}y_{p}\right\|-\left(t_{p}-1\right)\left\|y_{p}\right\|$
$>\left(\frac{1}{t}-\frac{1}{p}\right)t_{p}+\left\|t_{p}x_{p}-y_{p}\right\|+1-t_{p}$
$\geq\left(\frac{1}{t}-\frac{1}{p}\right)t_{p}+t_{p}+1-t_{p}$
$=\left(\frac{1}{t}-\frac{1}{p}\right)t_{p}+1.$
@@
Suppose that $x,y\in S_{X}$ and $x\perp_{B}y$, then we have $\|x-y\|\geq 1$, By the definition of $\delta_{x}(1)$, we can obtain that $\|x+y\|\leq 2\left(1-\delta_{x}(1)\right)$.
@@
Note that a Banach space $X$ is uniformly non-square if and only if $J^{B}(X)<2$ [12], by the above Proposition 12, we deduce the following Corollary.
Corollary 1.
@@
This shows that $J^{B}(X)\geq 2$. It should be noted that the upper bound is 2, and combined with the conclusion that $J^{B}(X)=2$ if and only if the space $X$ is not uniformly nonsquare [6], as desired. ∎
@@ 4 Some applications of $A^{B}_{t}(X)$ and $D_{t}^{B}(X)$
Recall that an orthogonality relation $\perp$ is calls symmetric if the condition $x\perp y$ necessarily entails $\perp x$ . It is straightforward to verify that the canonical orthogonality defined on Hilbert spaces is symmetric. In contrast, it is a well-documented fact that Birkhoff orthogonality lacks this symmetry property. Nevertheless, James established the following conclusion in the work [24].
@@
[24] A Banach space $X$ with dimension $\operatorname{dim}X\geq 3$ is a Hilbert space if and only if Birkhoff orthogonality is symmetric on $X$.
@@
[25] A two-dimensional Banach space in the Birkhoff orthogonality is symmetric, defined as a Radon plane.
From Proposition 3 and the fact that $X$ is a Radon plane, we obtain $A_{2}(X,B)\leq\frac{3}{2},$ which yields the following conclusion.
Corollary 2.
p+qp+qppq-qpq-p-qp-pqq
Figure 3: The unit sphere of the affine regular hexagon
By an affine-regular hexagon, any non-degenerate affine image obtained from a regular hexagon via affine transformation [26]. We observe that the space $\mathbb{R}^{2}$ with the norm $\ell_{\infty}-\ell_{1}$ is a Radon plane [24] (see Figure 2) and its unit sphere is an affine-regular hexagon. we have below example to achieve the upper bound of the corollary.
Example 3.
Then $A^{B}_{t}(X)=1+t-\frac{1}{2}\min\{1,t\}.$
Proof.
Suppose that $x=(1,0)$ and $y=(0,1)$. We can infer that $x,y\in S_{X}$ such that $x\perp_{B}y$. In addition, when $t\geq 1$, we have $\|x+ty\|=1$, $\|tx-y\|=t+1.$ Thus, it follows from Corollary 2 that we derive $A^{B}_{t}(X)=1+\frac{t}{2}$.
Similary, when $0\leq t\leq 1$, we have $\|x+ty\|=t$, $\|tx-y\|=t+1,$ so we can obtain that $A^{B}_{t}(X)=\frac{1}{2}+t$.
Thus we can complete the example.
Theorem 13.
Proof.
When $0\leq t\leq 1$, suppose that an affine-regular hexagon is $S_{X}$ , we have distinct points $p,q\in S_{X}$ such that $\pm p,\pm(p+q),\pm q$ are the vertices of $S_{X}$ (see Figure 3). Take $x=p+q$ and $y=p.$ Then it is evident that $x,y\in S_{X}$ with $x\perp_{B}y$.
In addition, we are also able to obtain
$\|tx-y\|=\|tq-(1-t)p\|=\|tq+(1-t)(-p)\|=1.$
Hence, by the definition of $A^{B}_{t}(X)$, we have
$1+\frac{t}{2}=\frac{\|x+ty\|+\|tx-y\|}{2}\leq A^{B}_{t}(X)\leq 1+\frac{t}{2},$
which means that $A^{B}_{t}(X)=1+\frac{t}{2}$.
Similary, when $t\geq 1$, we can deduce that $A^{B}_{t}(X)=t+\frac{1}{2}$. This completes the proof. ∎
@@
Remark 5.
In this paper, our geometric constants are only considered for the case involving a single parameter $t$. For the two-parameter case, similar arguments apply, and we can similarly establish their upper and lower bounds. Most of the other results remain valid as well.
At the end of the article, we would like to pose some open questions:
Problem 1:
Due to the properties of Birkhoff orthogonality, it is an interesting topic to incorporate the classical von Neumann-Jordan constant into the framework of Birkhoff orthogonality and redefine it as follows:
$C_{\mathrm{NJ}}(X)=\sup\left\{\frac{\|x+y\|^{2}+\|x-y\|^{2}}{2\left(\|x\|^{2}+\|y\|^{2}\right)}:x,y\in X,\ (x,y)\neq(0,0),\ x\perp_{B}y\right\}.$
Of course, the estimation of related inequalities also presents considerable difficulties. Furthermore, it is also natural to incorporate the $p$-th von Neumann-Jordan constant [39] into Birkhoff orthogonality. For related results on the $p$-th Birkhoff orthogonality, we refer the reader to [40].
Problem 2:
Remark 3 motivates us to investigate whether we can improve the proof of the well-known Lemma 6 by defining the constant as follows: set
$F(X)=\sup_{t>0}\left\{\frac{\|x+ty\|-\|tx+y\|}{2}:x,y\in S_{X}\right\}.$